# Evaluating ML Robustness in GNSS Interference Classification, Characterization & Localization

### v2

@@ Abstract
Jamming devices disrupt signals from the global navigation satellite system (GNSS) and pose a significant threat, as they compromise the robustness of accurate positioning. The detection of anomalies within frequency snapshots is crucial to counteract these interferences effectively. A critical preliminary countermeasure involves the reliable classification of interferences and the characterization and localization of jamming devices. This paper introduces an extensive dataset comprising snapshots obtained from a low-frequency antenna that capture various generated interferences within a large-scale environment, including controlled multipath effects. Our objective is to assess the resilience of machine learning (ML) models against environmental changes, such as multipath effects, variations in interference attributes, such as interference class, bandwidth, and signal power, the accuracy jamming device localization, and the constraints imposed by snapshot input lengths. Furthermore, we evaluate the performance of a diverse set of 129 distinct vision encoder models across all tasks. By analyzing the aleatoric and epistemic uncertainties, we demonstrate the adaptability of our model in generalizing across diverse facets, thus establishing its suitability for real-world applications. Dataset: https://gitlab.cc-asp.fraunhofer.de/darcy_gnss/controlled_low_frequency
@@
Assessing the robustness of a model stands as a critical aspect, typically necessitating its application across multiple independent datasets. Given the constraints associated with collecting extensive datasets, attention is often directed towards analyzing the distributions wherein the algorithm exhibits suboptimal performance [13, 14]. Notably, shifts in datasets between training and testing sets commonly lead to a decline in model efficacy, prompting efforts to alleviate such distributional discrepancies [15, 16, 17]. One viable approach involves evaluating either the latent space or the predictive uncertainty [18]. However, model robustness is often assessed on synthetic datasets, leaving uncertainties regarding the model’s resilience in the face of distributional shifts observed in authentic data scenarios [19]. Numerous real-world datasets exhibit such distributional shifts, evident in domains like handwriting recognition [20] and EEG signal classification [18], where the origins of data shifts are multifaceted.
@@ II Related Work
GNSS Interference Classification. Classical approaches utilize specific features, including automatic gain control (AGC) [23], received signal strength (RSS) values combined with time difference of arrival techniques [24], particle filters [25], and conformal seven-element antenna arrays [26], as well as methods designed for low-earth orbit applications [27]. In contrast, several recent ML-based approaches, such as convolutional neural networks (CNNs), support vector machines (SVMs), and random forests, have been proposed [28, 29, 30, 31, 32, 33]. Ott et al. [3] proposed an uncertainty-based Quadruplet loss aiming a more continuous representation between positive and negative interference pairs. They utilize a dataset resembling a snapshot-based real-world dataset, featuring similar interference classes, however, with varying sampling rates. Raichur et al. [4] used the same dataset to adapt to novel interference classes through continual learning. Jdidi et al. [2] proposed an unsupervised method of adapting to diverse environment-specific factors, i.e., multipath effects, dynamics, and variations in signal strength. Raichur et al. [1] introduced a crowdsourcing method utilizing smartphone-based features to localize the source of any detected interference. Brieger et al. [12] integrated both the spatial and temporal relationships between samples using a joint loss function and a late fusion technique. Their dataset was acquired in a controlled indoor environment. As ResNet18 [34] proved to be robust for interference classification, we also employ ResNet18 for feature extraction.
@@ III-B Definition of Experiments
In this section, we articulate the experiments feasible with the proposed datasets, aimed at assessing the model’s robustness. These experiments encompass: (1) Classification of seven distinct categories, comprising six types of interference and a single category denoting non-interference. (2) Characterization of interference by categorizing the energy level; the signal power with noise. (3) Characterization of interference by categorizing the BW. (4) Evaluation of diverse train-test splits: a dependent split, where both the training and testing sets possess identical interference characteristics (i.e., BW), and an independent split, where such characteristics differ. This facilitates the assessment of model generalizability. (5) Classification of antenna regions to anticipate the angle of interference. (6) Utilization of the recorded dataset for interference localization, entailing prediction of the signal generator’s position. Accordingly, the model is trained to predict positions within the open hall (16 positions), on the gallery (30 positions), or a combination of both (46 positions). (7) Cross-validation across 12 distinct training and testing scenarios, enabling analysis of environmental variability and robustness against multipath effects, particularly increased absorption. (8) Investigation of snapshot length, initially set at a maximum of 34, which is systematically reduced to a minimum of 1. This exploration aims to determine the minimal data requirement for achieving high classification accuracies across each task.
@@ V Evaluation
| Intf. type (dep.) | Signal power | 94.44 7.99 | 93.92 9.00 | - | 0.81 dB |
@@
Hardware and Training Setup. For all experiments, we use Nvidia Tesla V100-SXM2 GPUs with 32 GB VRAM equipped with Core Xeon CPUs and 192 GB RAM. We use the vanilla SGD optimizer with a multi-step learning rate of , decay of , and train the models for 200 epochs. We train each model 10 times and present the mean and standard variance accuracy and weighted F2-score for classification tasks, as well as the mean absolute error (MAE) for regression tasks. Table III provides a summary of all evaluation results.
@@
Signal Power. The model yields an accuracy of 86.42% on the signal power classification task, yet it exhibits misclassifications between signal power 6 dB and 8 dB, as well as between and (see Figure 4). However, detecting no interference remains reliable. The model’s capability to accurately distinguish between various signal powers is hindered by an increase of aleatoric and epistemic uncertainty. An alternative normalization procedure could potentially amplify the distinctions among snapshots. The regression task poses a challenge for the model, resulting in an error of . Although the classification accuracy marginally reduces to 96.15% in an independent task, the accuracy on the signal power drops significantly to 55.96% (refer to Table III), due to the intricate generalization toward unfamiliar signal powers.
@@
Architecture Benchmark. Figure 8 presents the evaluation results of 19 tsai models for all interference characterization tasks. The plot includes shaded regions representing standard deviations, which highlight the uncertainty or variability in the models’ predictions. ResNet and ResCNN, along with smaller models such as FCN, GRU-FCN, and InceptionTime, achieve high classification accuracies for both interference and signal power. However, ResNet consistently outperforms all other models across all tasks. Figure 7 benchmarks 110 Hugging Face models, revealing varying accuracy trends, with some models maintaining consistently high performance while others exhibit significant fluctuations. The best-performing models include ResNet variants, EfficientNet, and ConvNeXt, which demonstrate strong and stable accuracy, making them the most reliable choices for interference characterization and related tasks.

### v3

@@ Abstract
Jamming devices disrupt signals from the global navigation satellite system (GNSS) and pose a significant threat, as they compromise the robustness of accurate positioning. The detection of anomalies within frequency snapshots is crucial to counteract these interferences effectively. A critical preliminary countermeasure involves the reliable classification of interferences and the characterization and localization of jamming devices. This paper introduces an extensive dataset comprising snapshots obtained from a low-frequency antenna that capture various generated interferences within a large-scale environment, including controlled multipath effects. Our objective is to assess the resilience of machine learning (ML) models against environmental changes, such as multipath effects, variations in interference attributes, such as interference class, bandwidth, and signal power, the accuracy of jamming device localization, and the constraints imposed by snapshot input lengths. Furthermore, we evaluate the performance of a diverse set of 129 distinct vision encoder models across all tasks. By analyzing the aleatoric and epistemic uncertainties, we demonstrate the adaptability of our model in generalizing across diverse facets, thus establishing its suitability for real-world applications. Dataset: https://gitlab.cc-asp.fraunhofer.de/darcy_gnss/controlled_low_frequency
@@
Assessing the robustness of a model stands as a critical aspect, typically necessitating its application across multiple independent datasets. Given the constraints associated with collecting extensive datasets, attention is often directed towards analyzing the distributions wherein the algorithm exhibits suboptimal performance [13, 14]. Notably, shifts in datasets between training and testing sets commonly lead to a decline in model efficacy, prompting efforts to alleviate such distributional discrepancies [15, 16, 17, 18]. One viable approach involves evaluating either the latent space or the predictive uncertainty [19]. However, model robustness is often assessed on synthetic datasets, leaving uncertainties regarding the model’s resilience in the face of distributional shifts observed in authentic data scenarios [20]. Numerous real-world datasets exhibit such distributional shifts, evident in domains like handwriting recognition [21] and EEG signal classification [19], where the origins of data shifts are multifaceted.
@@ II Related Work
GNSS Interference Classification. Classical approaches utilize specific features, including automatic gain control (AGC) [24], received signal strength (RSS) values combined with time difference of arrival techniques [25], particle filters [26], and conformal seven-element antenna arrays [27], as well as methods designed for low-earth orbit applications [28]. In contrast, several recent ML-based approaches, such as convolutional neural networks (CNNs), support vector machines (SVMs), and random forests, have been proposed [29, 30, 31, 32, 33, 34]. Ott et al. [3, 35] proposed an uncertainty-based Quadruplet loss aiming a more continuous representation between positive and negative interference pairs. They utilize a dataset resembling a snapshot-based real-world dataset, featuring similar interference classes, however, with varying sampling rates. Raichur et al. [4] used the same dataset to adapt to novel interference classes through continual learning. Jdidi et al. [2] proposed an unsupervised method of adapting to diverse environment-specific factors, i.e., multipath effects, dynamics, and variations in signal strength. Raichur et al. [1] introduced a crowdsourcing method utilizing smartphone-based features to localize the source of any detected interference. Brieger et al. [12] integrated both the spatial and temporal relationships between samples using a joint loss function and a late fusion technique. Their dataset was acquired in a controlled indoor environment. As ResNet18 [36] proved to be robust for interference classification, we also employ ResNet18 for feature extraction.
@@ III-B Definition of Experiments
In this section, we articulate the experiments feasible with the proposed datasets, aimed at assessing the model’s robustness. These experiments encompass: (1) Classification of seven distinct categories, comprising six types of interference and a single category denoting non-interference. (2) Characterization of interference by categorizing the energy level; the signal power with noise. (3) Characterization of interference by categorizing the BW. (4) Evaluation of diverse train-test splits: a dependent split, where both the training and testing sets possess identical interference characteristics (i.e., BW), and an independent split, where such characteristics differ. This facilitates the assessment of model generalizability. (5) Classification of antenna regions to anticipate the angle of interference. (6) Utilization of the recorded dataset for interference localization, entailing prediction of the signal generator’s position (relative position between the MXG and the antenna based on labels). Accordingly, the model is trained to predict positions within the open hall (16 positions), on the gallery (30 positions), or a combination of both (46 positions). (7) Cross-validation across 12 distinct training and testing scenarios, enabling analysis of environmental variability and robustness against multipath effects, particularly increased absorption. (8) Investigation of snapshot length, initially set at a maximum of 34, which is systematically reduced to a minimum of 1. This exploration aims to determine the minimal data requirement for achieving high classification accuracies across each task.
@@ V Evaluation
| Intf. type (dep.) | Signal power | 94.44 7.99 | 93.92 9.00 | - | 0.81 dBm |
@@
For all experiments, we utilize the same training setup.111We use Nvidia Tesla V100-SXM2 GPUs with 32 GB VRAM equipped with Core Xeon CPUs and 192 GB RAM, the vanilla SGD optimizer with a multi-step learning rate of , decay of , and train the models for 200 epochs. We train each model 10 times and present the mean and standard variance accuracy (ratio between correctly classified samples and total number of samples) and weighted F2-score () for classification tasks, as well as the mean absolute error (MAE) for regression tasks. Table III provides a summary of all evaluation results.
@@
Signal Power. The model yields an accuracy of 86.42% on the signal power classification task, yet it exhibits misclassifications between signal power and , as well as between and (see Figure 4). However, detecting no interference remains reliable. The model’s capability to accurately distinguish between various signal powers is hindered by an increase of aleatoric and epistemic uncertainty. An alternative normalization procedure could potentially amplify the distinctions among snapshots. The regression task poses a challenge for the model, resulting in an error of . Although the classification accuracy marginally reduces to 96.15% in an independent task, the accuracy on the signal power drops significantly to 55.96% (refer to Table III), due to the intricate generalization toward unfamiliar signal powers.
@@
Architecture Benchmark. Figure 8 presents results of 19 tsai models for all characterization tasks. The plot includes shaded regions representing standard deviations, which highlight the uncertainty or variability in the models’ predictions. ResNet and ResCNN, along with smaller models such as FCN, GRU-FCN, and InceptionTime, achieve high classification accuracies for both interference and signal power. However, ResNet consistently outperforms all other models across all tasks. Figure 7 benchmarks 110 Hugging Face models, revealing varying accuracy trends, with some models maintaining consistently high performance while others exhibit significant fluctuations. The best-performing models include ResNet variants, EfficientNet, and ConvNeXt, which demonstrate strong and stable accuracy, making them the most reliable choices for interference characterization and related tasks.